26 May 2017
Friday 26th of May 2017

J.P. is blowing snow back on top of the orange balloon. After the snow has hardened, the balloon is deflated and pulled out leaving a cast snow pipe. A rope is attached to the balloon so that later a cable can be pulled through.

Third day with good weather in a row.

We have now for three days been blessed with good weather in camp, and in all fairness, I need to mention this after all my complaining about the weather in the last few weeks. It is still nice and cold; but the absence of strong wind makes a world of difference. It is a bit ironic though, that at this stage of camp work, most people work underground. Life in camp has found a new rhythm. With 28 in camp the main dome is crowded; but people seem to enjoy the company of each other, and with a little flexibility everyone has a nail for clothes and a shelf for private stuff. One of the draw backs of being many is the need for tools, and Nicolai, our carpenter, gave a speech at supper on how people could borrow tools; but they had to return them, otherwise he was left with an empty workshop.

What we did today:

  1. Made cable tunnel for heavy cable from main generator to drill trench.
  2. Drilling 7 runs.
  3. Logging ice cores.
  4. Working on infrastructure in science trench.
  5. Outfitting warm laboratories. Setting up isotope equipment and physical properties equipment.
  6. Testing radar.
  7. Trouble shooting the new freezer unit.
  8. Repairing snowblower. Blower spirals needed welding.
  9. German snow sampling.
  10. Japanese aerosol sampling and snow experiments.
  11. Water vapour sampling, methane sampling, boundary layer and meteorology station working.

Ad. 2: Drilllers report: Continued liquid drilling, 7 runs with HT electronics and RECAP drill, 1.7 m to 2.0 m cores. Drillers depth: 132.29m. The vacuum tank for suction of the mixture of drilling chips and drilling fluid now works as intended. In the vacuum tank, nicknamed “the pig”, heating elements warm up the mixture and the chips melt. The water and fluid then separates, and the water can be drained. The drilling fluid is recycled.

Ad.3: Logging depth, start: 132.96 m. Logging depth, end: 143.05 m.

Ad.7: For a couple of days, the freezer unit could not maintain a low temperature in the core extraction room. When running, the compressor sounded strange. It turned out that two phases had been reversed, thus making the compressor go backwards. The system appears to be running fine again.

Weather today: Sunshine. Few clouds in the afternoon. Temp. - 12°C to - 26°C. Wind: 5 - 10 kt from W turning SW. Visibility: To horizon.

FL, J.P. Steffensen

← Previous entry   Next entry →